OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (6, 1) and has slope 5/6. Please show all of your work.

OpenStudy (anonymous):

the slope is given so the equation can now be written as: y = 5/6x + b the b will be the y-intercept. since you are given a point (you are technically given and x and y-value) which you can plug into the equation and solve for b: 1 = 5/6(6) + b 1 = 5 + b 1-5 = b -4 = b so the equation now becomes: y = 5/6x -4

OpenStudy (anonymous):

yes (: